six years ago i was three times as old as my brother and now i am twice as old; find our present ages?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

let brother's age = x
my age = 2x
..
6 years ago our ages were:
brother - x-6
myself 2x-6
..
I was 3 time his age then
2x-6 = 3(x-6)
2x-6 = 3x-18
3x-2x=18-6
x=12 years brother's age
My age now = 24 years.
..
Your feedback is very valuable for improving the solution.

Answer by stanbon(75887)

duplicate
Cheers,
Stan H.